Ahmad has a far point of a mere 50 cm. He uses a contact lens to see the distant object clearly. What is/are the condition(s) about the contact lens?

A. The contact lens is a concave lens.

B. The contact lens is a diverging lens.

C. The focal length of the contact lens is -50 cm.

D. The power of the contact lens is -2.0 D


1
Expert's answer
2021-07-12T16:24:04-0400

Answer

Its Near sightedness because he can not see beyond 50cm.

So he has to use concave lens. Which is also known as diverging lens.


By lenses formula

"\frac{1}{f}=\frac{1}{v}-\frac{1}{u}"


V=-50cm

And "u=\infin"


So

"f=-50cm"


And

Power

"P=\frac{1}{f}=\frac{1}{-50}=-2.0D"


So A, B, C and D all are correct.